Three consecutive integers are such that their sum has the same value as twice the second integer subtracted from 20. What are the three integers?
Answer by mananth(16949)

let the integers be x ,x+1,x+2
x+x+1+x+2=20-2(x+1)
3x+3=20-2x-2
3x+2x=20-3-2
5x=15
/5
x=3
3,4,5 are the integers
...
m.ananth@hotmail.ca